OpenStudy (anonymous):

write the prime factorisation for 84

jimthompson5910 (jim_thompson5910):

84 = 42*2 42 = 21*2 21 = 7*3 So 84 = 2*2*3*7

OpenStudy (anonymous):

it is: 2*2*3*7

OpenStudy (anonymous):

2*2*3*7=81